what is the distributive property of
3x+2+5x

Answer by ptaylor(1894)

what is the distributive property of
3x+2+5x--------------------------it's 3x+5x or x(3+5)

The distributive law states that a(b+c)=ab+ac. Now lets look at your problem and re-write it:
3x+2+5x=x*3+x*5+2 or x(3+5)+2 or 8x+2

Hope this helps---ptaylor